I need some help with this.
24 / x - 3 = 72 / x + 3

Answer by benni1013(165)

Pick a denominator to multiply each side by.
I am choosing (x+3)
Multiply each side by x+3/1
You should have 24(x+3)/x-3=72==>24x+72/x-3=72
Multiply each side by (x-3)/1: 24x+72=72(x-3)==>24x+72=72x-216
Subtract the small coefficient from the larger: 72=48x-216
Add 216 to the other side: 288=48x
Divide each side by 48: 288/48=x==>x=6